OpenStudy (anonymous):

use elimination to solve the system. 3x+5y=6 2x-4y=4

OpenStudy (anonymous):

3x + 5y = 6 --> (2) 2x - 4y = 4 --> (-3) ---------------- 6x + 10y = 12 (result of multiplying by 2) -6x + 12y = -12 (result of multiplying by -3) ----------------add 22y = 0 y = 0 now sub 0 in for y in either of the above equations 3x + 5y = 6 3x + 5(0) = 6 3x = 6 x = 2 check.... 2x - 4y = 4 2(2) - 4(0) = 4 4 = 4 (correct) x = 2 and y = 0 or (2,0) any questions ?

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.